OpenStudy (anonymous):

how do i solve this im just checking my answer 4log 4u-4log4v

OpenStudy (mertsj):

What is your answer?

OpenStudy (anonymous):

log4(u^4/v^4)

OpenStudy (aum):

\[ \log_4\left (\frac uv \right)^4 \]

OpenStudy (aum):

My answer is the same as yours.

TheSmartOne (thesmartone):

except simplified :)

OpenStudy (jhannybean):

\[\large \sf n \log(x) \iff \log(x^n)\]

TheSmartOne (thesmartone):

\[\log_{a} b-\log_{a} c=\log_{a} (\frac{b}{c})\]